A local hamburger shop sold a combined total of 822 hamburgers and cheeseburgers on Wednesday. There were 72 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Wednesday?

Answer by suhasbhat(3)

Let h be hamburgers sold and c be the cheeseburgers sold.
h + c = 822. ---> eqn(1)
But c = h + 72. (because 72 more cheese burgers were sold than hamburgers).
Now substitute (h+72) for c in eqn(1).
h + (h+72) = 822.
2 h + 72 = 822.
2h = 822 - 72
2h = 750.
h = 375.
A total of 375 hamburgers and 375+72 cheeseburgers were sold on the given Wednesday.